# On the Robustness of RSMA to Adversarial BD-RIS-Induced Interference

### v2

@@ III-C Quantifying Robustness of RSMA
We propose two performance metrics to evaluate the robustness of RSMA against RIS-induced attacks, namely rate degradation and robustness index. The rate degradation quantifies the performance loss caused by an attack, computed as the difference between the baseline performance in a secure environment and the degraded performance during an attack. To this end, the common rate is assumed to be uniformly allocated among the users, i.e., $R^{c}/U$. The overall effective rate for user $u$ is then the sum of this portion of the common rate and the individual private rate $R^{p}_{u}$. The rate degradation experienced by the $u$th user is defined as
@@
The robustness index $\kappa$ complements the rate degradation metric by evaluating the proportion of baseline performance retained during an attack. This normalized metric quantifies the system’s ability to maintain performance under RIS-induced attacks relative to its baseline rates $\bar{R}^{p}_{u}$ and $\bar{R}^{c}$. The robustness index is defined as
@@ III-D CSI Acquisition
Under the assumption of reciprocity, the downlink channels can be estimated at the BS based on uplink measurements from pilot signals transmitted by the users. As explained, during this phase, the attacker can configure the BD-RIS to either absorb or randomly reflect incoming signals. In absorption mode, the pilot signals impinging on the BD-RIS are prevented from reaching the BS. In the random reflective mode, on the other hand, the pilot signals will characterize the BD-RIS-related propagation paths. However, the attacker reconfigures the reflection coefficients in the data transmission phase, making the acquired CSI inaccurate. Thus, in both modes, only imperfect estimates of the channels $\mathbf{h}_{u}$ are accessible to the BS. Adding to the BD-RIS distortions, other rapid changes in the propagation environment and hardware limitations can also introduce errors in the estimation process. As a result, the direct link estimates will be a combination of the true channels and error components, modeled as:
@@
As for the attacker, we assume that it can only estimate the channels associated with its BD-RIS, also modeled as imperfect. Specifically, the estimates of the BS-RIS channel matrix and the RIS-user channel vector are given by
@@ III-E BS Precoder Design
The private precoder should ensure that the interference observed at user $u$ due to transmissions intended for any other user $u^{\prime}$ is negligible, i.e., $\mathbf{h}_{u^{\prime}}^{H}\mathbf{w}_{u}^{p}\approx 0$, $\forall\,u^{\prime}\neq u$. This objective is achieved via a regularized zero-forcing approach, which effectively mitigates inter-user interference while balancing noise enhancement. Specifically, the BS first constructs the estimated direct-link channel matrix as
@@ III-F Power Allocation
A meaningful way to perform power allocation is to keep the interference resulting from imperfect CSI approximately at the same level as the noise floor [11]. By doing this, the degradation of users’ data rates can be mitigated. In our system, which considers imperfect SIC, the noise floor for the private messages is a combination of the thermal noise $\sigma^{2}$ and the residual SIC interference $\chi^{c}_{u}$. Therefore, we aim to ensure that the SINR expression for the private stream, $\gamma_{u}^{p}$, is on the same order as the sum of this effective noise power, i.e., $\sum_{i\neq u}|\mathbf{h}_{u}^{H}\mathbf{w}^{p}_{i}|^{2}P\alpha^{p}_{u}\approx\xi\left|\mathbf{h}_{u}^{H}\mathbf{w}^{c}\right|^{2}P\alpha^{c}+\sigma^{2}$. To achieve this goal, we first adopt a uniform power allocation across the private messages, i.e., $\alpha^{p}=\alpha^{p}_{u},\forall u\in\mathcal{U}$. Then, we can substitute $\alpha^{c}=1-U\alpha^{p}$ and show for the $u$th user that:
@@
Note that we need to compute an $\alpha^{p}$ capable of balancing the interference terms across all users in such a way that the system’s sum rate is maximized. Given this goal, we prioritize the user with the best channel conditions, i.e., the one that contributes the most to the sum rate. To this end, we select as reference the best performance indicator given by:
@@ IV-B Aligned interference attack
A more sophisticated adversarial strategy is to optimize the BD-RIS reflection coefficients to maximize the interference induced at the users. As demonstrated in [6] for conventional RIS systems, if the adversary can acquire partial RIS-associated CSI [^1] To acquire the necessary CSI, an adversary could exploit semi-passive RIS architectures, where active sensing elements are incorporated to estimate the RIS channels over time [31]., it becomes possible to strategically steer the reflected power toward the users. In this subsection, we generalize this strategy to BD-RIS architectures.
@@
where $\bm{u}_{ij}$ is a unit vector of length $\frac{D(D+1)}{2}$ with a single non-zero entry at position $(j-1)D+i-\frac{j(j-1)}{2}]$, and $\mathbf{T}_{ij}$ is a $D\times D$ matrix with ones at positions $(i,j)$ and $(j,i)$, and zeros elsewhere [32, Definitions 3.2a and 3.2b].
@@ 5   Compute: $\mathbf{S}_{i}\leftarrow(\mathbf{G}^{T}\otimes\mathbf{g}_{i}^{H})$;
$\mathbf{S}\leftarrow[\sqrt{\mu_{1}}\mathbf{S}_{1}^{H},\cdots,\sqrt{\mu_{U}}\mathbf{S}_{U}^{H}]^{H}.$
@@ 14   Compute the block-row matrix:
$\mathbf{J}_{i}\leftarrow\begin{bmatrix}\mathbf{S}_{i1}\mathbf{D}_{1}&\mathbf{S}_{i2}\mathbf{D}_{2}&\cdots&\mathbf{S}_{iG}\mathbf{D}_{G}\end{bmatrix}\,$.
@@
$\mathbf{J}\leftarrow\begin{bmatrix}\sqrt{\mu_{1}}\mathbf{J}_{1}^{T},\sqrt{\mu_{2}}\mathbf{J}_{2}^{T},\cdots,\sqrt{\mu_{U}}\mathbf{J}_{U}^{T}\end{bmatrix}^{T}\,.$
@@
20   Extract reduced vector: $\bm{\theta}_{g}\leftarrow[\bm{\varphi}]_{\mathcal{R}_{g}}$, where $\mathcal{R}_{g}\subset\left\{1,\cdots,\sum_{g\in\mathcal{G}}\frac{D_{g}(D_{g}+1)}{2}\right\}$ is the index set for group $g$, with $|\mathcal{R}_{g}|=\frac{D_{g}(D_{g}+1)}{2}$.
@@
Given that $\sum_{i\in\mathcal{U}}\mathbf{S}_{i}^{H}\mathbf{S}_{i}=\left[\mathbf{S}_{1}^{H},\cdots,\mathbf{S}_{U}^{H}\right]^{H}\left[\mathbf{S}_{1}^{H},\cdots,\mathbf{S}_{U}^{H}\right]$, we can define the tall weighted interference matrix: $\mathbf{S}\triangleq[\sqrt{\mu_{1}}\mathbf{S}_{1}^{H},\cdots,\sqrt{\mu_{U}}\mathbf{S}_{U}^{H}]^{H}\in\mathbb{C}^{UN\times D^{2}}$ and rewrite the expression in (IV-B) as the convex quadratic objective $\bm{\theta}^{H}\mathbf{D}^{T}\mathbf{S}^{H}\mathbf{S}\mathbf{D}\bm{\theta}$. This vectorized reformulation automatically satisfies the symmetry constraint in (28b) by construction. On the other hand, the non-convex unitary constraint in (28c) is relaxed with a convex $\ell_{2}$-norm bound $\|\bm{\theta}\|_{2}\leq 1$, which allow us to formulate a tractable optimization problem:
@@
The aligned interference attack is now generalized for the group-connected architecture. Let us consider a BD-RIS partitioned into $G$ independent groups, i.e., $\bm{\Theta}=\mathrm{bdiag}(\bm{\Theta}_{1},\dots,\bm{\Theta}_{G})$, with each block satisfying $\bm{\Theta}_{g}=\bm{\Theta}_{g}^{T}$ and $\bm{\Theta}_{g}^{H}\bm{\Theta}_{g}=\mathbf{I}$. Then, the optimized adversary attack strategy can be formulated as:
@@
Next, following Property IV-B, we construct a duplication matrix $\mathbf{D}_{g}\in\mathbb{R}^{D_{g}^{2}\times\frac{D_{g}(D_{g}+1)}{2}}$ to vectorize the symmetric matrix as $\mathrm{vec}(\bm{\Theta}_{g})=\mathbf{D}_{g}\bm{\theta}_{g}$ for each group, where $\bm{\theta}_{g}=\mathrm{vech}(\bm{\Theta}_{g})\in\mathbb{C}^{\frac{D_{g}(D_{g}+1)}{2}}$ denotes the $g$th reduced coefficient vector. We also define $\mathbf{S}_{ig}=(\mathbf{G}_{g}^{T}\otimes\mathbf{g}_{ig}^{H})\in\mathbb{C}^{M\times D_{g}^{2}}$, and represent the set of RIS groups by $\mathcal{G}=\{1,\cdots,G\}$. Then, we rewrite the adversary objective as follows:
@@
$\bm{\varphi}\triangleq[\bm{\theta}_{1}^{T},\bm{\theta}_{2}^{T},\cdots,\bm{\theta}_{G}^{T}]^{T}\in\mathbb{C}^{\sum_{g\in\mathcal{G}}\frac{D_{g}(D_{g}+1)}{2}},$ (40)
@@
$\mathbf{J}_{i}\triangleq\begin{bmatrix}\mathbf{S}_{i1}\mathbf{D}_{1},\mathbf{S}_{i2}\mathbf{D}_{2},\cdots,\mathbf{S}_{iG}\mathbf{D}_{G}\end{bmatrix}$. (41)
@@
$\mathbf{J}\triangleq\begin{bmatrix}\sqrt{\mu_{1}}\mathbf{J}_{1}^{T},\sqrt{\mu_{2}}\mathbf{J}_{2}^{T},\cdots,\sqrt{\mu_{U}}\mathbf{J}_{U}^{T}\end{bmatrix}^{T}.$ (43)
@@
The solution to (44) can be obtained with the aid of the SVD, similarly to the problem (33). More specifically, we can decompose $\mathbf{J}=\mathbf{U}_{\mathbf{J}}\bm{\Sigma}_{\mathbf{J}}\mathbf{V}_{\mathbf{J}}^{H}$ and achieve the desired solution by computing $\bm{\varphi}=\mathbf{v}_{\mathbf{J},1}$, where $\mathbf{v}_{\mathbf{J},1}$ is the dominant right singular vector of $\mathbf{J}$, i.e., the first column of $\mathbf{V}_{\mathbf{J}}$. Finally, by letting $\mathcal{R}_{g}\subset\left\{1,\cdots,\sum_{g\in\mathcal{G}}\frac{D_{g}(D_{g}+1)}{2}\right\}$ denote the index subset corresponding to the $g$th reduced coefficient vector, with $|\mathcal{R}_{g}|=\frac{D_{g}(D_{g}+1)}{2}$, the relaxed coefficients for each independent RIS group can be obtained as $\bm{\theta}_{g}=[\bm{\varphi}]_{\mathcal{R}_{g}}$, and $\bm{\Phi}_{g}=\mathrm{unvec}(\mathbf{D}_{g}\bm{\theta}_{g})$, $\forall g\in\mathcal{G}$. Recalling Property IV-A, we compute the Takagi factorization $\bm{\Phi}_{g}=\mathbf{U}_{\bm{\Phi}_{g}}\bm{\Sigma}_{\bm{\Phi}_{g}}\mathbf{U}_{\bm{\Phi}_{g}}^{T}$ and project each $\bm{\Phi}_{g}$ onto its closest unitary matrix, as follows:
@@
Finally, the full BD-RIS reflection matrix is constructed as $\bm{\Theta}^{*}=\mathrm{bdiag}(\bm{\Theta}^{*}_{1},\cdots,\bm{\Theta}^{*}_{G})$.

### v3

@@ III-C Quantifying Robustness of RSMA
We propose two performance metrics to evaluate the robustness of RSMA against RIS-induced attacks, namely rate degradation and robustness index. The rate degradation quantifies the performance loss caused by an attack, defined as the difference between the baseline performance in a secure environment and the degraded performance during an attack. To this end, the common rate is assumed to be uniformly allocated among the users, i.e., $R^{c}/U$. The overall effective rate for user $u$ is then the sum of this portion of the common rate and the individual private rate $R^{p}_{u}$. The rate degradation experienced by the $u$th user is defined as
@@
The robustness index $\kappa$ complements the rate degradation metric by quantifying the proportion of baseline performance retained during an attack. This normalized metric quantifies the system’s ability to maintain performance under RIS-induced attacks relative to its baseline rates $\bar{R}^{p}_{u}$ and $\bar{R}^{c}$. The robustness index is defined as
@@ III-D CSI Acquisition
Under the assumption of reciprocity, the downlink channels can be estimated at the BS based on uplink measurements from pilot signals transmitted by the users. As explained, during this phase, the attacker can configure the BD-RIS to either absorb or randomly reflect incoming signals. In absorption mode, the pilot signals impinging on the BD-RIS are prevented from reaching the BS. In the random reflective mode, on the other hand, the pilot signals will characterize the BD-RIS-related propagation paths. However, the attacker reconfigures the reflection coefficients during the data transmission phase, rendering the acquired CSI inaccurate. Thus, in both modes, only imperfect estimates of the channels $\mathbf{h}_{u}$ are accessible to the BS. Adding to the BD-RIS distortions, other rapid changes in the propagation environment and hardware limitations can also introduce errors in the estimation process. As a result, the direct link estimates will be a combination of the true channels and error components, modeled as:
@@
As for the attacker, we assume it can only estimate the channels associated with its BD-RIS and that these estimates are imperfect. Specifically, the estimates of the BS-RIS channel matrix and the RIS-user channel vector are given by
@@
It is noteworthy that channel acquisition is a well-known practical challenge in RIS-aided systems, due to limited sensing hardware, calibration and synchronization challenges, or insufficient observation time due to mobility, and even more so when the RIS operates in a stand-alone manner without direct coordination with the BS, as would be the case for an attacker. One potential solution is to exploit architectures with embedded sensing capabilities, such as the concept of a semi-passive RIS proposed in [23], where a small subset of elements is equipped with dedicated sensing circuitry to acquire partial CSI and reconstruct the remaining channels at the RIS controller by passively sensing existing pilot transmissions. A related approach is the hybrid reflecting-and-sensing RIS [51], in which elements are dual-purpose, simultaneously reflecting and sensing a portion of the impinging signal to enable channel estimation from existing pilots. In a realistic adversarial deployment, however, such acquisition remains challenging, resulting in imperfect estimates, as we assume in our model.
@@ III-E BS Precoder Design
The private precoder should ensure that the interference observed at user $u$ due to transmissions intended for any other user $u^{\prime}$ is negligible, i.e., $\mathbf{h}_{u^{\prime}}^{H}\mathbf{w}_{u}^{p}\approx 0$, $\forall\,u^{\prime}\neq u$. This objective is achieved via a regularized zero-forcing approach [^3] Other precoding strategies could also be considered. For instance, iterative RSMA precoding strategies that jointly optimize transmit precoders and power allocation are expected to achieve higher absolute sum rates, e.g., WMMSE (WMMSE)-based designs [24], at the expense of a greater computational complexity. Incorporating WMMSE and other strategies into the proposed framework and investigating the impact on RSMA robustness is an interesting direction for future work., which effectively mitigates inter-user interference while balancing noise enhancement. Specifically, the BS first constructs the estimated direct-link channel matrix as
@@ III-F Power Allocation
A meaningful way to perform power allocation is to keep the interference resulting from imperfect CSI approximately at the same level as the noise floor [4]. By doing this, the degradation of users’ data rates can be mitigated. In our system, which considers imperfect SIC, the noise floor for the private messages is a linear combination of the thermal noise $\sigma^{2}$ and the residual SIC interference $\chi^{c}_{u}$. Therefore, we aim to ensure that the SINR expression for the private stream, $\gamma_{u}^{p}$, is on the same order as the sum of this effective noise power, i.e., $\sum_{i\neq u}|\mathbf{h}_{u}^{H}\mathbf{w}^{p}_{i}|^{2}P\alpha^{p}_{u}\approx\xi\left|\mathbf{h}_{u}^{H}\mathbf{w}^{c}\right|^{2}P\alpha^{c}+\sigma^{2}$. To achieve this goal, we first adopt a uniform power allocation across the private messages [^4] Allowing a more general joint power and precoder optimization across private and common streams is expected to achieve higher performance. However, due to the inherent complexity of such a joint design, it is left for future work., i.e., $\alpha^{p}=\alpha^{p}_{u},\forall u\in\mathcal{U}$. Then, we can substitute $\alpha^{c}=1-U\alpha^{p}$ and show for the $u$th user that:
@@
Note that we need to compute an $\alpha^{p}$ that balances the interference terms across all users so as to maximize the system’s sum rate. Given this goal, we prioritize the user with the best channel conditions, i.e., the one that contributes the most to the sum rate. To this end, we select as a reference the best performance indicator given by:
@@
This power allocation policy also provides intuition for the robustness trends observed in Section V. Since the BS is unaware of any BD-RIS manipulation, it performs precoding using the available corrupted CSI. The attack is then reflected in the effective SINR through increased residual inter-user interference due to beamforming mismatch, i.e., larger leakage terms $\sum_{i\neq u}\!\left|\mathbf{h}_{u}^{H}\mathbf{w}_{i}^{p}\right|^{2}$ in the denominator of the private SINR in (7). In particular, the proposed policy explicitly ties the private-stream power coefficient to the ratio in (22), where the denominator includes the private-stream leakage power $\sum_{i\neq u}\!\left|\mathbf{h}_{u}^{H}\mathbf{w}_{i}^{p}\right|^{2}P$. Thus, when the attack increases the leakage term, the resulting $\alpha_{p}$ decreases and the remaining power is shifted to the common stream, i.e., $\alpha_{c}=1-U\alpha_{p}$. As the transmit power increases, this mechanism becomes more pronounced because the interference terms dominate over $\sigma^{2}$, which promotes the allocation of more power to the common message. Therefore, even without attack awareness or explicit countermeasures, the RSMA adaptation mechanism intended to cope with standard imperfect CSI also mitigates the BD-RIS-induced degradation.
@@ IV-B Aligned interference attack
A more sophisticated adversarial strategy is to optimize the BD-RIS reflection coefficients to maximize the interference induced at the users. As demonstrated in [8] for conventional RIS systems, if the adversary can acquire partial RIS-associated CSI, it becomes possible to strategically steer the reflected power toward the users. In this subsection, we generalize this strategy to BD-RIS architectures.
@@
where $\bm{u}_{ij}$ is a unit vector of length $\frac{D(D+1)}{2}$ with a single non-zero entry at position $(j-1)D+i-\frac{j(j-1)}{2}]$, and $\mathbf{T}_{ij}$ is a $D\times D$ matrix with ones at positions $(i,j)$ and $(j,i)$, and zeros elsewhere [38, Definitions 3.2a and 3.2b].
@@ 5   Compute: $\mathbf{S}_{i}\leftarrow(\mathbf{G}^{T}\otimes\mathbf{g}_{i}^{H})$;
$\mathbf{S}\leftarrow[\sqrt{\mu_{1}}\mathbf{S}_{1}^{H},\ldots,\sqrt{\mu_{U}}\mathbf{S}_{U}^{H}]^{H}.$
@@ 14   Compute the block-row matrix:
$\mathbf{J}_{i}\leftarrow\begin{bmatrix}\mathbf{S}_{i1}\mathbf{D}_{1}&\mathbf{S}_{i2}\mathbf{D}_{2}&\ldots&\mathbf{S}_{iG}\mathbf{D}_{G}\end{bmatrix}\,$.
@@
$\mathbf{J}\leftarrow\begin{bmatrix}\sqrt{\mu_{1}}\mathbf{J}_{1}^{T},\sqrt{\mu_{2}}\mathbf{J}_{2}^{T},\ldots,\sqrt{\mu_{U}}\mathbf{J}_{U}^{T}\end{bmatrix}^{T}\,.$
@@
20   Extract reduced vector: $\bm{\theta}_{g}\leftarrow[\bm{\varphi}]_{\mathcal{R}_{g}}$, where $\mathcal{R}_{g}\subset\left\{1,\ldots,\sum_{g\in\mathcal{G}}\frac{D_{g}(D_{g}+1)}{2}\right\}$ is the index set for group $g$, with $|\mathcal{R}_{g}|=\frac{D_{g}(D_{g}+1)}{2}$.
@@
Given that $\sum_{i\in\mathcal{U}}\mathbf{S}_{i}^{H}\mathbf{S}_{i}=\left[\mathbf{S}_{1}^{H},\ldots,\mathbf{S}_{U}^{H}\right]^{H}\left[\mathbf{S}_{1}^{H},\cdots,\mathbf{S}_{U}^{H}\right]$, we can define the tall weighted interference matrix: $\mathbf{S}\triangleq[\sqrt{\mu_{1}}\mathbf{S}_{1}^{H},\ldots,\sqrt{\mu_{U}}\mathbf{S}_{U}^{H}]^{H}\in\mathbb{C}^{UN\times D^{2}}$ and rewrite the expression in (IV-B) as the convex quadratic objective $\bm{\theta}^{H}\mathbf{D}^{T}\mathbf{S}^{H}\mathbf{S}\mathbf{D}\bm{\theta}$. This vectorized reformulation automatically satisfies the symmetry constraint in (28b) by construction. On the other hand, the non-convex unitary constraint in (28c) is relaxed with a convex $\ell_{2}$-norm bound $\|\bm{\theta}\|_{2}\leq 1$, which allows us to formulate a tractable optimization problem:
@@
Note that relaxing the unitary constraint in (33) provides a tractable surrogate formulation of (28), whose solution captures the dominant interference direction of the original problem via SVD. Subsequently, the projection in (35) returns the closest symmetric unitary matrix in the Frobenius norm sense, as per Lemma IV-A, to the relaxed solution, and thus enforces feasibility but does not guarantee global optimality. Since (28) is inherently non-convex and its global optimum is generally intractable to characterize, determining the exact performance gap between the proposed suboptimal adversarial approach and the original formulation remains an open challenge.
@@
Herein, the aligned interference attack is specialized to the group-connected architecture. Let us consider a BD-RIS partitioned into $G$ independent groups, i.e., $\bm{\Theta}=\mathrm{bdiag}(\bm{\Theta}_{1},\dots,\bm{\Theta}_{G})$, with each block satisfying $\bm{\Theta}_{g}=\bm{\Theta}_{g}^{T}$ and $\bm{\Theta}_{g}^{H}\bm{\Theta}_{g}=\mathbf{I}$. Then, the optimized adversary attack strategy can be formulated as:
@@
Next, following Property IV-B, we construct a duplication matrix $\mathbf{D}_{g}\in\mathbb{R}^{D_{g}^{2}\times\frac{D_{g}(D_{g}+1)}{2}}$ to vectorize the symmetric matrix as $\mathrm{vec}(\bm{\Theta}_{g})=\mathbf{D}_{g}\bm{\theta}_{g}$ for each group, where $\bm{\theta}_{g}=\mathrm{vech}(\bm{\Theta}_{g})\in\mathbb{C}^{\frac{D_{g}(D_{g}+1)}{2}}$ denotes the $g$th reduced coefficient vector. We also define $\mathbf{S}_{ig}=(\mathbf{G}_{g}^{T}\otimes\mathbf{g}_{ig}^{H})\in\mathbb{C}^{M\times D_{g}^{2}}$, and represent the set of RIS groups by $\mathcal{G}=\{1,\ldots,G\}$. Then, we rewrite the adversary objective as follows:
@@
$\bm{\varphi}\triangleq[\bm{\theta}_{1}^{T},\bm{\theta}_{2}^{T},\ldots,\bm{\theta}_{G}^{T}]^{T}\in\mathbb{C}^{\sum_{g\in\mathcal{G}}\frac{D_{g}(D_{g}+1)}{2}},$ (40)
@@
$\mathbf{J}_{i}\triangleq\begin{bmatrix}\mathbf{S}_{i1}\mathbf{D}_{1},\mathbf{S}_{i2}\mathbf{D}_{2},\ldots,\mathbf{S}_{iG}\mathbf{D}_{G}\end{bmatrix}$. (41)
@@
$\mathbf{J}\triangleq\begin{bmatrix}\sqrt{\mu_{1}}\mathbf{J}_{1}^{T},\sqrt{\mu_{2}}\mathbf{J}_{2}^{T},\ldots,\sqrt{\mu_{U}}\mathbf{J}_{U}^{T}\end{bmatrix}^{T}.$ (43)
@@
TABLE I: Runtime (seconds) of optimized attacks for different RIS architectures
Number of RIS elements $D$
50 100 150 200 250 300
Fully Connected RIS (proposed, Algorithm 1) 0.20198 0.49007 0.89702 1.14418 1.84436 2.36110
Group-Connected RIS (proposed, Algorithm 2) 0.04219 0.05327 0.09056 0.10905 0.11455 0.13659
Single-Connected RIS (iterative baseline [8]) 0.70900 1.11176 1.36038 1.52885 1.82989 2.08470
The solution to (44) can be obtained with the aid of the SVD, similarly to the problem (33). More specifically, we can decompose $\mathbf{J}=\mathbf{U}_{\mathbf{J}}\bm{\Sigma}_{\mathbf{J}}\mathbf{V}_{\mathbf{J}}^{H}$ and achieve the desired solution by computing $\bm{\varphi}=\mathbf{v}_{\mathbf{J},1}$, where $\mathbf{v}_{\mathbf{J},1}$ is the dominant right singular vector of $\mathbf{J}$, i.e., the first column of $\mathbf{V}_{\mathbf{J}}$. Finally, by letting $\mathcal{R}_{g}\subset\left\{1,\ldots,\sum_{g\in\mathcal{G}}\frac{D_{g}(D_{g}+1)}{2}\right\}$ denote the index subset corresponding to the $g$th reduced coefficient vector, with $|\mathcal{R}_{g}|=\frac{D_{g}(D_{g}+1)}{2}$, the relaxed coefficients for each independent RIS group can be obtained as $\bm{\theta}_{g}=[\bm{\varphi}]_{\mathcal{R}_{g}}$, and $\bm{\Phi}_{g}=\mathrm{unvec}(\mathbf{D}_{g}\bm{\theta}_{g})$, $\forall g\in\mathcal{G}$. Recalling Property IV-A, we compute the Takagi factorization $\bm{\Phi}_{g}=\mathbf{U}_{\bm{\Phi}_{g}}\bm{\Sigma}_{\bm{\Phi}_{g}}\mathbf{U}_{\bm{\Phi}_{g}}^{T}$ and project each $\bm{\Phi}_{g}$ onto its closest unitary matrix, as follows:
@@
Finally, the full BD-RIS reflection matrix is constructed as $\bm{\Theta}^{*}=\mathrm{bdiag}(\bm{\Theta}^{*}_{1},\ldots,\bm{\Theta}^{*}_{G})$.